# In this paper, we investigate the mixed norm estimates for the operator $T$associated with a dilated plane curve $(ut,u\gamma(t))$, defined by

### v2

@@ 2. $\blacksquare$
Notice that $\phi(z,\cdot)$ is a homogeneous function of degree 111, which leads to that $\partial_{u}\phi(z,\cdot)$ is also a homogeneous function of degree 111 and $|\partial_{u}\phi(z,\cdot)|\lesssim|\xi|$. Then, the proof of the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate for $\partial_{u}(\mathcal{F}P_{k}f)$ is very similar as that for $\mathcal{F}P_{k}f$, but need to multiply a extra factor $2^{k}$. Therefore, by Lemma 3.6 with $r=q$ and Lemma 3.3, we can say
@@
On the other hand, since $F_{j,k}$ is a Fourier integral operator of order $-\frac{1}{2}$ satisfying the special assumption (2.10), Proposition 3.7 implies
@@
Here $\epsilon\in(0,\infty)$ is an arbitrary constant. Moreover, from [34], the estimates (3.1), (3.5), and (3.7) represent the endpoint estimates for $F_{j,k}$ when taking the $L^{\infty}$-norm in $u$. i.e., the points (0,0)(0,0)( 0 , 0 ), $(\frac{1}{2},\frac{1}{2})$, and $(\frac{2}{5},\frac{1}{5})$ are the endpoints. Consequently, $s_{-\frac{1}{2},0}(p,q,\infty)$ is the infimum such that (3.23) holds. Combining this with (3.22), we obtain $\frac{1}{r}\geq\max\{\frac{1}{p}-\frac{2}{q},\frac{3}{2p}-\frac{1}{2q}-\frac{1}{2},\frac{2}{p}-1\}$. This further implies
@@ 4.1.1 The Case $1\leq p\leq q\leq r\leq\infty$
Observe that $F_{j,k}P_{k}f=F_{j,k}f$, and $F_{j,k}$ is a Fourier integral operator of order $-\frac{1}{2}$ satisfying (2.10). By Lemma 3.3, we have
@@ 4.1.3 The Case $1\leq r<p\leq q\leq\infty$
$\tilde{s}_{C,\epsilon}(p,q,r)=\begin{cases}-\frac{2}{q}+\epsilon,&\text{for }q\geq 3p^{\prime}\text{ and }\frac{1}{p}\leq\frac{1}{4};\\ \frac{2}{p}-\frac{2}{q}-\frac{1}{2}+\epsilon,&\text{for }q\geq 3p^{\prime}\text{ and }\frac{1}{4}<\frac{1}{p}<\frac{1}{2}-\frac{1}{q};\\ \frac{1}{p}-\frac{3}{q}+\epsilon,&\text{for }q\geq 3p^{\prime}\text{ and }\frac{1}{p}\geq\frac{1}{2}-\frac{1}{q};\\ \frac{3}{2p}-\frac{3}{2q}-\frac{1}{2}+\epsilon,&\text{for }p^{\prime}<q<3p^{\prime};\\ \frac{2}{p}-\frac{1}{q}-1+\epsilon,&\text{for }q\leq p^{\prime}.\end{cases}$
If $\tilde{s}_{C,0}(p,q,r)<0$, then (4.8) holds with a decay in $k$, where $(\frac{1}{p},\frac{1}{q})$ belongs to the region with vertices (0,0)(0,0)( 0 , 0 ), $(\frac{1}{4},0)$, $(\frac{3}{8},\frac{1}{8})$, $(\frac{1}{2},\frac{1}{6})$, $(\frac{2}{3},\frac{1}{3})$, and (1,1)(1,1)( 1 , 1 ). Furthermore, the point $(\frac{3}{8},\frac{1}{8})$ can be removed by interpolation between $(\frac{1}{4},0)$ and $(\frac{1}{2},\frac{1}{6})$. Thus, (4.8) holds with some $s<0$ if $(\frac{1}{p},\frac{1}{q})$ satisfies
@@ 4.2 Necessary Part
From (4.3), setting $\bar{s}_{2,-\frac{1}{2}}(p,q)<0$ yields a decay estimate $F_{j,k}:L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$, where $(\frac{1}{p},\frac{1}{q})$ lies in the open region with vertices (0,0)(0,0)( 0 , 0 ), $(\frac{2}{3},\frac{1}{3})$, and (1,1)(1,1)( 1 , 1 ). Then, $T:L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$ holds on the intersection of this region with $1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)>0$. Thus, $T:L_{x}^{p}(\mathbb{R}^{2})\to L_{x}^{q}(\mathbb{R}^{2})$ holds on this region uniformly in $u\in[1,2]$. Combining this with [31, Theorem 2.1], we conclude that (0,0)(0,0)( 0 , 0 ), $(\frac{2}{3},\frac{1}{3})$, and (1,1)(1,1)( 1 , 1 ) are the endpoints of the estimate $F_{j,k}:L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$. Therefore, $\bar{s}_{2,-\frac{1}{2}}(p,q)$ is the infimum ensuring (4.3) holds.
@@
From (4.2), setting $s_{-\frac{1}{2},0}(p,q,q)<0$ yields the decay estimate $F_{j,k}:L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$ when $(\frac{1}{p},\frac{1}{q})$ lies in the open region with vertices (0,0)(0,0)( 0 , 0 ), $(\frac{1}{2},\frac{1}{6})$, $(\frac{2}{3},\frac{1}{3})$, and (1,1)(1,1)( 1 , 1 ). This result is nearly sharp by [31]. Similarly, if

### v3

@@ 3.2 The necessity of the region of $(\frac{1}{p},\frac{1}{q})$ for all $r\in[1,\infty]$
Notice that $\phi(z,\cdot)$ is a homogeneous function of degree $1$, which leads to that $\partial_{u}\phi(z,\cdot)$ is also a homogeneous function of degree $1$ and $|\partial_{u}\phi(z,\cdot)|\lesssim|\xi|$. Then, the proof of the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate for $\partial_{u}(\mathcal{F}P_{k}f)$ is very similar as that for $\mathcal{F}P_{k}f$, but need to multiply a extra factor $2^{k}$. Therefore, by Lemma 3.2 with $r=q$ and Lemma 3.4, we can say
@@
On the other hand, since $F_{j,k}$ is a Fourier integral operator of order $\frac{1}{2}$ satisfying the special assumption (2.10), Proposition 3.7 implies
@@
Here $\epsilon\in(0,\infty)$ is an arbitrary constant. Moreover, from [35], the estimates (3.1), (3.5), and (3.7) represent the endpoint estimates for $F_{j,k}$ when taking the $L^{\infty}$-norm in $u$. i.e., the points $(0,0)$, $(\frac{1}{2},\frac{1}{2})$, and $(\frac{2}{5},\frac{1}{5})$ are the endpoints. Consequently, $s_{-\frac{1}{2},0}(p,q,\infty)$ is the infimum such that (3.26) holds. Combining this with (3.25), we obtain $\frac{1}{r}\geq\max\{\frac{1}{p}-\frac{2}{q},\frac{3}{2p}-\frac{1}{2q}-\frac{1}{2},\frac{2}{p}-1\}$. This further implies
@@ 4.1.1 The Case $1\leq p\leq q\leq r\leq\infty$
Observe that $F_{j,k}P_{k}f=F_{j,k}f$, and $F_{j,k}$ is a Fourier integral operator of order $\frac{1}{2}$ satisfying (2.10). By Lemma 3.4, we have
@@ 4.1.3 The Case $1\leq r<p\leq q\leq\infty$
[…]
If $\tilde{s}_{C,0}(p,q,r)<0$, then (4.8) holds with a decay in $k$, where $(\frac{1}{p},\frac{1}{q})$ belongs to the region with vertices $(0,0)$, $(\frac{1}{4},0)$, $(\frac{3}{8},\frac{1}{8})$, $(\frac{1}{2},\frac{1}{6})$, $(\frac{2}{3},\frac{1}{3})$, and $(1,1)$. Furthermore, the point $(\frac{3}{8},\frac{1}{8})$ can be removed by interpolation between $(\frac{1}{4},0)$ and $(\frac{1}{2},\frac{1}{6})$. Thus, (4.8) holds with some $s<0$ if $(\frac{1}{p},\frac{1}{q})$ satisfies
@@ 4.2 Necessary Part
From (4.3), setting $\bar{s}_{2,-\frac{1}{2}}(p,q)<0$ yields a decay estimate $F_{j,k}:L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$, where $(\frac{1}{p},\frac{1}{q})$ lies in the open region with vertices $(0,0)$, $(\frac{2}{3},\frac{1}{3})$, and $(1,1)$. Then, $T:L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$ holds on the intersection of this region with $1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)>0$. Thus, $T:L_{x}^{p}(\mathbb{R}^{2})\to L_{x}^{q}(\mathbb{R}^{2})$ holds on this region uniformly in $u\in[1,2]$. Combining this with [31, Theorem 2.1], we conclude that $(0,0)$, $(\frac{2}{3},\frac{1}{3})$, and $(1,1)$ are the endpoints of the estimate $F_{j,k}:L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$. Therefore, $\bar{s}_{2,-\frac{1}{2}}(p,q)$ is the infimum ensuring (4.3) holds.
@@
From (4.2), setting $s_{-\frac{1}{2},0}(p,q,q)<0$ yields the decay estimate $F_{j,k}:L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$ when $(\frac{1}{p},\frac{1}{q})$ lies in the open region with vertices $(0,0)$, $(\frac{1}{2},\frac{1}{6})$, $(\frac{2}{3},\frac{1}{3})$, and $(1,1)$. This result is nearly sharp by [31]. Similarly, if